OpenStudy (anonymous):

how to find value of tan(5pi/2)?

OpenStudy (anonymous):

Imagine where is 5pi/2 located in a unit circle. I think it's in the positive y exis, coterminal with pi/2 or 90 degrees. tangent is sine/cosine, but cos(5pi/2) = 0 so tangent (5pi/2) is undefined

OpenStudy (dumbcow):

let 2x = 5pi/2 x = 5pi/4 use double angle identities \[\tan(5\pi/2) = \frac{2\tan(5\pi/4)}{1-\tan^{2}(5\pi/4)}\]

OpenStudy (dumbcow):

haha yeah i did it the long way essentially 5pi/2 is same as pi/2 tan(pi/2) is undefined

OpenStudy (anonymous):

thank u all

OpenStudy (anonymous):

Just curious, what happens when tan of a number is und? What happens to the answer?

OpenStudy (anonymous):

As in this case?

OpenStudy (anonymous):

so what about 7pi/8? what is its value bcause it is not coterminal?

OpenStudy (dumbcow):

it will go to either pos or neg infinity depending on which side you are approaching so limit DNE, thus there is no answer its undefined

OpenStudy (anonymous):

The y value of the function increases without bound (or decreases toward negative infinity).

OpenStudy (anonymous):

if you check out the graph of tangent function, it has an asymptote at pi/2, 3pi/2 and other coterminal angles.

OpenStudy (anonymous):

@dumbcow so the final answer simply is undefined?

OpenStudy (anonymous):

graph never reach the undefined points of tangent function

OpenStudy (dumbcow):

yes or more specifically, 5pi/2 is not in the domain of tan x

OpenStudy (anonymous):

I was shown to do it like this 5pi/2-1(2pi), I am not getting the right answer, what am I doing wrong?

OpenStudy (anonymous):

The question is, Find the measurement in radians of the least positive angle that is coterminal with the given angle 5pi/2? Now it was shown to me to use this form 5pi/2-1(2pi) I need help?